Plan for 20170205 UT E-time

Local Mean Sidereal Time at midnight =   8 36 00.1
 Sunset ( 1000 m horizon):   17 59 MST; Sunrise:   7 07 MST
 Evening twilight:  19 18 MST;  LMST at evening twilight:   3 53
 Morning twilight:   5 48 MST;  LMST at morning twilight:  14 25

PROPOSED SEQUENCE (version of 03-Feb-2017)

TCS (First Half Hour = 2-2.5 UT, configuration=LUCI1/IRTC2)

  • DONE Test new TCS patches on-sky (JohnH, Michele)
    • DONE GCS hotspot code (priority=high, time=0.5 hr, seeing=any)
    • DONE DIMM patch (in parallel to other activities)

AO (Next 3 hours = 2.5-5.5 UT)

  • DONE Test new version 16.4 of AOS subsystem (see Issue 6563) (better error messages, time=included in next task)
  • DONE Test new AO Supervisor Code with FLAO (UAO built from local SVN as user AOEng) (Greg+Xianyu, priority=high, time=2 hours, seeing=<1.5, clouds=thin)
    • DONE Test normal AO observation including dithering. (many scripts exist)
    • DONE Test AO observation with recenter=1 (time=0.2 hours)
    • DONE Test AO observation created by OT (time=0.3 hours) (Michelle making script)
  • DONE Tip-Tilt measurement with IRTC2 (John+Doug time=0.5 hours, seeing<1.5, clouds=thin)

Science (5.5-8 UT including reconfiguration time)

  • AIP DDT X-ray Cluster (RADEC=09+61, anytime but first/last hour)
    • DONE MODS/MODS MOS spectra (Olga, time=2.2 hours, clouds<thin, seeing=OK)

LFBG Checkout and Alignment (Second Half)

  • AGW1 Re-alignment (John, priority=high, 1 hour, seeing<1)
    Procedure for measuring the alignment of the hotspot to the rotator center in BGTechProcedures
  • DONE AGW1 pupil wobble check (John+Doug, priority=high, 1 hour, seeing<1.5)
  • LUCI1 transform data (Olga+John, priority=high, 2 hours, seeing<1.2) (need LUCI scripts, or take data with IDL)
    (Have these LUCI scripts:Procedure for measuring the alignment of the hotspot to the rotator center
    	StoneE.xml	07:14   +00:14	
    	StoneF.xml	08:04   +00.30
                           StoneG 10:30
                           StoneH 11:22
                           StoneI  12:54
                           StoneJ 13:45
                           StoneK 15:41
                           StoneL 16:30 (disagrees with Dave's script coordinates)
    	StoneL.xml	18:59   +00:02
    	StoneM.xml	18:44   -00:12
    (parallel IRTC2 transform data, only possible with IDL)
  • AGW1 field aberration data (John, priority=med, 1.5 hour, seeing<1)


POSTPONE UNTIL LATER FOR LACK OF TIME

  • Observe Al's mag 10.8 asteroid (16 Psyche RADEC=1115+05) with LUCI1AO (cloudy backup with good seeing, can handle ~2 mags of uniform clouds) (priority=optional, time=0.5 hour)

  • Test NS guiding at 150 arcsec/hour (15 sec wfs exposure) (bad seeing and/or cloudy backup) (priority=optional, time=0.7 hr, seeing=OK)

  • Pointing Data (only when seeing not suitable for AGW and AO work)
    • Stable Temperature Binocular Set ~28 stars (2 hours, any seeing)
    • Cloudy Binocular Set (~56 stars, any seeing, up to 2 mag clouds)
    • Test existing new binocular pointing model (collimation models still missing from John)

  • ESM Observing tests (John+LUCIop has test scripts near RA=11)

  • LBTI checkout for Feb 5 observing (2 hours, TBD) (probably not needed since LBTI was operating Jan 29, and little new SW)

  • Test new UAO code with LBTI (depends on status of merging Alfio's fixes above - not ready)

  • MODS1 transform and rotator zeropoint (Olga, time=1.5 hours, seeing<1.5, clouds=thin)

* John's AO QSOs (real LUCI AO, not only ESM as above, seeing<1.3, few clouds)
(K-band for Q0751+34, 1 hour between 5-9 UT)
(H,K for Q1130+60, 2 hours after 6 UT)

POSTPONE UNTIL LATER FOR TECHNICAL REASONS

  • RFBG
    • AGW2 Re-alignment
    • AGW2 pupil wobble check
    • LUCI2 transform
    • AGW2 field aberrations
    • Check focus on LUCI2 wfs and WFS

-- %USERSIG{JohnHill - 2017-02-03}%
Topic revision: r6 - 05 Feb 2017, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback